OpenStudy (anonymous):

Two boys want to use a seesaw, and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what have you got so far?

OpenStudy (anonymous):

i have got the numbers all written down and the things they equal... I also have made the multiplication division by x1 * d1 over w2

jimthompson5910 (jim_thompson5910):

does it give you a diagram or something?

jimthompson5910 (jim_thompson5910):

the numbers are missing

jimthompson5910 (jim_thompson5910):

oh they just want you to solve for d2 only?

OpenStudy (anonymous):

how do i figure out the numbers? no it doesn't

OpenStudy (anonymous):

im in the literal equations lesson

jimthompson5910 (jim_thompson5910):

I think they just want you to solve for d2

jimthompson5910 (jim_thompson5910):

to do that, we simply divide both sides by w2 to isolate d2 \[\Large w_{1}*d_{1} = w_{2}*d_{2}\] \[\Large \frac{w_{1}*d_{1}}{w_{2}} = \frac{w_{2}*d_{2}}{w_{2}}\] \[\Large \frac{w_{1}*d_{1}}{w_{2}} = d_{2}\] \[\Large d_{2} = \frac{w_{1}*d_{1}}{w_{2}}\]

OpenStudy (anonymous):

yes thats what i believe well they say they want to see your rewrite formula for the solving of d2

jimthompson5910 (jim_thompson5910):

yeah basically get d2 on one side all by itself

OpenStudy (anonymous):

then the basic formula is meaning isolate the variable sort of like d=r*t

jimthompson5910 (jim_thompson5910):

yeah d is isolated

OpenStudy (anonymous):

ok so the general formula is ?

jimthompson5910 (jim_thompson5910):

I just wrote it above

jimthompson5910 (jim_thompson5910):

when I solved for d2

OpenStudy (anonymous):

i got it thanks :D